# Universal Physics Transformers: A Framework For Efficiently Scaling Neural Operators

### v2

@@ B.3 ShapeNet-Car
We test on the dataset generated by Umetani & Bickel (2018), which consists 889 car shapes from ShapeNet (Chang et al., 2015) where side mirrors, spoilers and tires were removed. We randomly split the 889 simulations into 700 train and 189 test samples. Each sample consists of 3682 mesh points, including a small amount of points that are not part of the car mesh. We filter all points that do not belong to the car mesh, resulting in 3586 points per sample. For each point, we predict the associated pressure value. As there is no notion of time and the inflow velocity is constant, no feature modulation is necessary. We use the transformer positional encoding (Vaswani et al., 2017), and, therefore rescale all positions to the range [0, 200].
@@
For GINO, we create a positional embedding of mesh- and grid-positions which are then used as input to a 3 layer MLP with GELU (Hendrycks & Gimpel, 2016) non-linearities to create messages. Messages from mesh-positions within a radius of each grid-position are accumulated which serves as input to the FNO part of GINO. The FNO uses 64 hidden dimensions with 16/24/32 modes per dimension for grid resolutions $32^{3}/48^{3}/64^{3}$, respectively. The GINO decoder encodes the query positions and again uses a 3 layer MLP with GELU non-linearities to create messages. Each query position aggregates the messages from grid-points within a radius. The radius for message aggregation is set to $10$. When using SDF features, the SDF features are encoded via a 2 layer MLP to the same dimension as the hidden dimension, and added onto the positional encoding of each grid point.
@@
When additionally using SDF features as input, the SDF features are encoded by a shallow ConvNeXt V2 (Woo et al., 2023) that processes the SDF features into 512 (8x8x8) tokens, which are concatenated to the perceiver tokens. To distinguish between the two token types, a learnable vector per type is added to each of the tokens.
@@
Extended results.
Table 4 extends the main paper results with the additional resolution of $48^{3}$.
Profiling memory and runtime.
We evaluate the memory per sample and runtime per epoch (Table 2 and Table 4) by searching the largest possible batchsize via a binary search. To get the memory consumption per sample, we divide the peak memory consumption by the largest possible batchsize. For the runtime per epoch, we conduct a short benchmark run with the largest possible batchsize take and extrapolate the time to 1 epoch. All benchmarks are conducted on a A100 80GB SXM GPU.
Model SDF #Tokens MSE Mem. [GB]
U-Net 0 $64^{3}$ 6.13 1.3
FNO 0 $64^{3}$ 4.04 3.8
GINO 0 $64^{3}$ 2.34 19.8
UPT (ours) 0 64 2.31 0.6
U-Net 32 $32^{3}$ 3.66 0.2
FNO 32 $32^{3}$ 3.31 0.5
GINO 32 $32^{3}$ 2.90 2.1
UPT (ours) 32 $10248^{3}+$ 1024 2.35 2.7
U-Net 48 $48^{3}$ 3.33 0.5
FNO 48 $48^{3}$ 3.29 1.6
GINO 48 $48^{3}$ 2.57 7.9
UPT (ours) 48 $10248^{3}+$ 1024 2.25 2.7
U-Net 64 $64^{3}$ 2.83 1.3
FNO 64 $64^{3}$ 3.26 3.8
GINO 64 $64^{3}$ 2.14 19.8
UPT (ours) 64 $10248^{3}+$ 1024 2.24 2.7
Table 4: Normalized test MSE for ShapeNet-Car pressure prediction. The loss is multiplied by 100. Memory denotes the amount required for a forward and backward pass of a single sample. UPTs can model the dynamics with a fraction of latent tokens compared to other models.
B.4 Transient flows
@@
The dataset consists of 10K simulations which we split into 8K training simulations, 1K validation simulations and 1K test simulations. Each simulations has a length of 100 seconds with a $\Delta t$ of 1s. The trajectories are randomly generated 2D windtunnel simulations with 1-4 objects (circles of varying size) placed randomly in the tunnel.
The dataset is converted to float16 precision to save storage. From all training samples, data statistics are extracted to normalize inputs to approximately mean 0 and standard deviation 1. In order to get robust statistics, the normalization parameters are calculated only from values within the inter-quartile range. As the Pytorch function to calculate quartiles is not supported to handle such a large amount of values, we assume that values follow a Gaussian distribution with mean and variance from the data, which allows us to infer the inter-quartile range from the Gaussian CDF.
@@
To efficiently use multi-GPU setups, we sample the same number of mesh points for each GPU. Without this modification, the FLOPS in the encoder would fluctuate drastically per GPU, which results in “busy waiting” time for GPUs with less FLOPS in the encoder. We also keep the number of mesh points constant between batches to avoid memory fluctuations, which allows using a larger batchsize. We rescale the x positions from the range $[-0.5,0.5]$ to $[0,200]$ and the y positions from $[-0.5,1]$ to $[0,300]$. We do this to avoid complications with the positional embedding (Vaswani et al., 2017) which was designed for positive integer positions. To create a graph out of mesh points, we use a radius graph with $r=5$ and limit the maximum number of graph edges per node to 32 to avoid memory fluctuations. We choose $r=5$ as it covers the whole domain when using a grid resolution of $64^{2}$.
@@
We train all models for 100 epochs using a batchsize of 1024 where we use gradient accumulation if the full batchsize does not fit into memory. Following common practices, we use AdamW (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) for U-Net, FNO and GINO. As we encountered training instabilities when training UPTs on this dataset, we change the optimizer for UPTs to Lion (Chen et al., 2023). These instabilities manifest in sudden loss spikes from which the model can not recover. Using the Lion optimizer together with float32 precision training solves these instabilities in our experiments. When training with Lion, we use a learning rate of $5\times 10^{-5}$ and $1\times 10^{-4}$ otherwise. We also use float32 for GINO as we found that this improves performance. U-Net and FNO do not benefit from float32 and are therefore trained in bfloat16 precision.
Other hyperparameters are chosen based on common practices when training Vision Transformers (ViT) Dosovitskiy et al. (2021). We linearly increase the learning rate for the first 10 epochs Loshchilov & Hutter (2017) followed by a cosine decay. A weight decay of $0.05$ is used when training with AdamW, which is increased to 0.5 when using Lion as suggested in Chen et al. (2023)). Architectural choices like the number of attention heads, the latent dimension for each attention head or the expansion dimension for the transformer’s MLP are copied from ViTs.
We find float32 precision to be beneficial for GINO and UPTs, which is likely due to positional embeddings being too inaccurate with bfloat16 precision. Training with float16 instead of bfloat16 (to preserve the precision of the positional embedding while speeding up training) resulted in NaN values due to overflows. Therefore, we train all GINO and UPT models with float32. We found this to be very important as otherwise UPTs become unstable over the course of training, leading to a large loss spike from which the model can not recover. During model development, we tried various ways to solve this instabilities where using the Lion optimizer was one of them. With fp32 training, UPTs work similarly well with AdamW but we kept Lion as we already trained models and re-training all models is quite expensive.
@@
B.5 Lagrangian fluid mechanics
Datasets
We use the Taylor-Green vortex datasets from Toshev et al. (2023b) for our Lagrangian experiments, comprising two dimensions (TGV2D) and three dimensions (TGV3D). The Taylor-Green vortex problem is characterized by a distinctive initial velocity field without an external driving force, resulting in a static decrease in kinetic energy over time. The TGV2D dataset includes 2500 particles, a sequence length of 126, and is characterized by a Reynolds number (Re) of 100. The data is split into 100/50/50 for training, validation, and testing, respectively. The TGV3D dataset consists of 8000 particles, a sequence length of 61, and a Reynolds number of 50. The data is partitioned into 200/100/100 for training, validation, and testing, respectively.
Baselines
The Graph Network-based Simulator (GNS) model (Sanchez-Gonzalez et al., 2020) is a popular learned surrogate for physical particle-based simulations. The architecture is kept simple, adhering to the encoder-processor-decoder principle (Battaglia et al., 2018), where the processor consists of multiple graph network blocks (Battaglia et al., 2016). The Steerable E(3)-equivariant Graph Neural Network (SEGNN) architecture (Brandstetter et al., 2022c) is a general implementation of an E($3$) equivariant GNN, where layers are directly conditioned on steerable attributes for both nodes and edges. The main building blocks are steerable MLPs, i.e., stacks of learnable linear Clebsch-Gordan tensor products interleaved with gated non-linearities (Weiler et al., 2018). SEGNN layers are message-passing layers (Gilmer et al., 2017) where steerable MLPs replace the traditional non-equivariant MLPs for both message and node update functions.
We utilize the checkpoints provided by Toshev et al. (2023b) for the comparisons in 4.3. Both GNS and SEGNN baseline models comprise 10 layers with latent dimensions of 128 and 64, respectively. The maximum irreps order of SEGNN are $l=1$, for more details see Toshev et al. (2023b).
Hyperparameters
We train using a batchsize of 128 for 50 epochs with a warmup phase of 10 epochs. We sample inputs between $50\%$ to $100\%$ of total particles. For optimization, we use AdamW (Loshchilov & Hutter, 2019) for all experiments with learning rate $10^{-3}$ and weight decay of $0.05$.
UPT architecture
For the TGV2D and TGV3D experiments, we use $n_{\text{s}}=256$ and $n_{\text{s}}=512$ supernodes, respectively. For TGV3D more supernodes are required due to the increased number of particles. Message passing is done with a maximum of four input points per supernode.
B.6 Out-of-distribution generalization study
@@
Generalization to more objects.
@@
Generalization to higher velocities.
@@
The in-distribution dataset contains random velocities sampled from $v\in[0.01,0.06]$ m/s. The velocity is used (together with the timestep) as input to the modulation layers (e.g. DiT Peebles & Xie (2023) modulation) which modulate the features of all layers in the network. This leads to OOD velocities distorting all features within a forward pass. The center plot of Fig. 12 shows that UPT has the best OOD generalization among all models. Notably, the performance of GINO drastically drops for higher velocities.
Generalization to different geometries.
@@
The in-distribution dataset contains randomly placed circles of varying size where the mesh is generated via an adaptive meshing algorithm. To investigate robustness to the meshing algorithm, we generate a OOD dataset with a uniform meshing algorithm. In this dataset, mesh cells are approximately uniform, i.eṫhe distance between two points is roughly the same. This is in contrast to an adaptive mesh, where regions around an object have more mesh cells and therefore the distance between two points is smaller in these regions. Additionally, we investigate generalization to different obstacle geometries by using polygons (with up to 9 edges) or triangles instead of circles. Note that even though polygons are a more ”complicated” shape, they are more reminiscent of a circle than triangles. The size of the obstacle is also varied here. For simplicity, the number of objects per simulation is set to 1 in this study. Note that U-Net and FNO interpolate the mesh onto a 2D grid and therefore their distribution of input positions does not change here, only the dynamics of the simulation. For GINO and UPT, also the distribution of input position changes. The right side of Fig. 12 shows that UPT achieves the best performances on OOD meshes, despite having to adjust to a different input position distribution.
@@
Figure 12: OOD generalization study. Trained 68M parameter models are evaluated on OOD datasets with more objects, higher inflow velocities and different geometries. ”ID” refers to an adaptive meshing algorithm with circles as obstacles. ”UMesh” changes the meshing algorithm from adaptive to uniform. ”Triangle” uses triangles instead of circles and ”Polygon” uses polygon obstacles (with up to 9 edges) instead of circles. UPTs have similar OOD generalization capabilities as other models, outperforming them in all evaluations. Grey indicates OOD settings.
B.7 Qualitative study of scaling limits details

### v3

@@ D.4 ShapeNet-Car
[…]
@@
For GINO, we create a positional embedding of mesh- and grid-positions which are then used as input to a 3 layer MLP with GELU [37] non-linearities to create messages. Messages from mesh-positions within a radius of each grid-position are accumulated which serves as input to the FNO part of GINO. The FNO uses 64 hidden dimensions with 16/24/32 modes per dimension for grid resolutions $32^{3}/48^{3}/64^{3}$, respectively. The GINO decoder encodes the query positions and again uses a 3 layer MLP with GELU non-linearities to create messages. Each query position aggregates the messages from grid-points within a radius. The radius for message aggregation is set to $10$. When using SDF features, the SDF features are encoded via a 2 layer MLP to the same dimension as the hidden dimension, and added onto the positional encoding of each grid point.
@@
When additionally using SDF features as input, the SDF features are encoded by a shallow ConvNeXt V2 [113] that processes the SDF features into 512 (8x8x8) tokens, which are concatenated to the perceiver tokens. To distinguish between the two token types, a learnable vector per type is added to each of the tokens.
@@
Extended results
Table 4 gives a detailed overview of all results from ShapeNet-Car with varying resolutions and comparison to baseline models.
Table 4: Normalized test MSE for ShapeNet-Car pressure prediction. Memory is the amount of memory required to train on a single sample. UPTs can model the dynamics with a fraction of latent tokens compared to other models. SDF additionally uses the signed distance function from each gridpoint to the geometry as input features. To include SDF features into UPT, we encode the SDF features with resolution $32^{3}$, $48^{3}$ or $64^{3}$ into $8^{3}$ tokens using a shallow ConvNeXt V2 [113] and concatenate these tokens to the tokens coming from the UPT encoder. To balance the number of SDF tokens with the number of latent tokens, we increase the number of latent tokens to 1024 in the settings where we use SDF features for UPT. Runtime is measured on an A100 GPU.
Model SDF #Tokens MSE [1e-2] Mem. [GB] Runtime per Epoch [s]
U-Net $64^{3}$ 6.13 1.3 86
FNO $64^{3}$ 4.04 3.8 148
GINO $64^{3}$ 2.34 19.8 900
UPT (ours) 64 2.31 0.6 4
U-Net $32^{3}$ 3.66 0.2 11
FNO $32^{3}$ 3.31 0.5 17
GINO $32^{3}$ 2.90 2.1 103
UPT (ours) $10248^{3}+$ 1024 2.35 2.7 156
U-Net $48^{3}$ 3.33 0.5 35
FNO $48^{3}$ 3.29 1.6 64
GINO $48^{3}$ 2.57 7.9 360
UPT (ours) $10248^{3}+$ 1024 2.25 2.7 156
U-Net $64^{3}$ 2.83 1.3 86
FNO $64^{3}$ 3.26 3.8 148
GINO $64^{3}$ 2.14 19.8 900
UPT (ours) $10248^{3}+$ 1024 2.24 2.7 156
Profiling memory and runtime
We evaluate the memory per sample and runtime per epoch (Table 4) by searching the largest possible batchsize via a binary search. To get the memory consumption per sample, we divide the peak memory consumption by the largest possible batchsize. For the runtime per epoch, we conduct a short benchmark run with the largest possible batchsize take and extrapolate the time to 1 epoch. All benchmarks are conducted on a A100 80GB SXM GPU.
D.5 Transient flows
@@
The dataset consists of 10K simulations which we split into 8K training simulations, 1K validation simulations and 1K test simulations. Each simulations has a length of 100 seconds where a $\Delta t$ of 1s is used to train neural surrogate models. Note that this $\Delta t$ is different from the $\Delta t$ used by the numerical solver, which requires a much lower $\Delta t$ to remain stable. Concretely, we set the temporal update of the numerical solver initially to $\Delta t=0.05$s. If instabilities occur, the trajectory is rerun with smaller $\Delta t$. Overall, each trajectory comprises 2K timesteps at the coarsest $\Delta t$ setting and 200K at the finest ($\Delta t=0.0005$s). The trajectories are randomly generated 2D windtunnel simulations with 1-4 objects (circles of varying size) placed randomly in the tunnel. The uni-directional inflow velocity varies between 0.01 to 0.06 m/s. Each simulation contains between 29K and 59K mesh points where each point has three features: pressure, and the $x-$ and $y-$component of the velocity. In total, the dataset is converted to float16 precision to save storage and amounts to roughly 235GB.
From all training samples, data statistics are extracted to normalize inputs to approximately mean 0 and standard deviation 1. In order to get robust statistics, the normalization parameters are calculated only from values within the inter-quartile range. As the Pytorch function to calculate quartiles is not supported to handle such a large amount of values, we assume that values follow a Gaussian distribution with mean and variance from the data, which allows us to infer the inter-quartile range from the Gaussian CDF.
@@
To efficiently use multi-GPU setups, we sample the same number of mesh points for each GPU. Without this modification, the FLOPS in the encoder would fluctuate drastically per GPU, which results in “busy waiting” time for GPUs with less FLOPS in the encoder. We also keep the number of mesh points constant between batches to avoid memory fluctuations, which allows using a larger batchsize. We rescale the x positions from the range $[-0.5,0.5]$ to $[0,200]$ and the y positions from $[-0.5,1]$ to $[0,300]$. We do this to avoid complications with the positional embedding [106] which was designed for positive integer positions. To create a graph out of mesh points, we use a radius graph with $r=5$ and limit the maximum number of graph edges per node to 32 to avoid memory fluctuations. We choose $r=5$ as it covers the whole domain when using a grid resolution of $64^{2}$.
@@
We train all models for 100 epochs using a batchsize of 1024 where we use gradient accumulation if the full batchsize does not fit into memory. Following common practices, we use AdamW [41, 59] for U-Net, FNO and GINO. As we encountered training instabilities when training UPTs on this dataset, we change the optimizer for UPTs to Lion [21]. These instabilities manifest in sudden loss spikes from which the model can not recover. Using the Lion optimizer together with float32 precision training solves these instabilities in our experiments. When training with Lion, we use a learning rate of $5\times 10^{-5}$ and $1\times 10^{-4}$ otherwise. We also use float32 for GINO as we found that this improves performance. U-Net and FNO do not benefit from float32 and are therefore trained in bfloat16 precision.
Other hyperparameters are chosen based on common practices when training ViTs [26]. We linearly increase the learning rate for the first 10 epochs [58] followed by a cosine decay. A weight decay of $0.05$ is used when training with AdamW, which is increased to 0.5 when using Lion as suggested in [21]. Architectural choices like the number of attention heads, the latent dimension for each attention head or the expansion dimension for the transformer’s MLP are copied from ViTs.
We find float32 precision to be beneficial for GINO and UPTs, which is likely due to positional embeddings being too inaccurate with bfloat16 precision. Training with float16 instead of bfloat16 (to increase the precision of the positional embedding while speeding up training) resulted in NaN values due to overflows. Therefore, we train all GINO and UPT models with float32. We found this to be very important as otherwise UPTs become unstable over the course of training, leading to a large loss spike from which the model can not recover. During model development, we tried various ways to solve this instabilities where using the Lion optimizer was one of them. With fp32 training, UPTs work similarly well with AdamW but we kept Lion as we already trained models and re-training all models is quite expensive.
@@ D.5.1 Visualization of rollouts
D.5.2 Supernode discretization convergence
While one would ideally use lots of supernodes also during training, it also requires more computational power. Therefore, we investigate increasing the number of supernodes only during inference. We use the 68M parameter models from Sec. 4.2 and increase the number of supernodes during inference. Figure 8 shows that UPTs are discretization convergent and inference performance can be improved by using more supernodes for inference than during training.
Figure 8: We study discretization convergence by varying the number the number of gridpoints/supernodes of models that were trained 2K supernodes (UPT) or 4K grid points (GINO). UPT demonstrates a stable performance across different number of supernodes even though it has never seen that number of supernodes during training. Increasing the number of supernodes even improves the performance of UPT slightly. In contrast, the performance of GINO plummets when the number of gridpoints is different during inference.
D.5.3 Impact of a larger latent space
D.5.4 Performance of smaller models
Figure 5 could suggest that UPT does not scale well with parameter size due to the steeper decline of GINO. However, the scaling of GINO only looks good because GINO underperforms in contrast to UPT (GINO-68M is worse than UPT-8M). For well trained (UPT) models it gets increasingly difficult to improve the loss further. Ideally, one would show this by training larger GINO models, however this is prohibitively expensive (68M parameter models already take 450 A100 hours per model). We therefore go in the other directions and train even smaller UPT models that achieve a similar loss to the GINO models and compare scaling there. In Figure 10 right, we compare UPT 1M/2M/8M against GINO 8M/17M/68M. UPT shows similar scaling on that loss scale.
We strongly hypothesis that the effect of larger UPT models would become apparent in even more challenging settings or even larger datasets. However, challenging large-scale datasets are hard to come by, which is why we created one ourselves. Creating even larger and more complex ones is beyond the scope of our work as it exceeds our current resource budget, but it is definitely an interesting direction for future research/applications.
Figure 10: UPT is much more expressive than GINO and shows good scaling when increasing model size. UPT-1M achieves comparable performance to GINO-17M using 17x less parameters.
D.5.5 Scalability with dataset size
We show the scalability and data efficiency of UPTs by training UPT-8M on subsets of the data used for the transient flow experiments (we train on 2K and 4K out of the 8K train simulations). The results in Figure 11 right show that UPT scales well with data and is data efficient, achieving comparable results to GINO-8M with 4x less data.
Figure 11: UPT scales well with more data and is data efficient, achieving comparable results to competitors with 4x less data.
D.5.6 Out-of-distribution generalization study
@@
Generalization to more objects
@@
Generalization to higher velocities
@@
The in-distribution dataset contains random velocities sampled from $v\in[0.01,0.06]$ m/s. The velocity is used (together with the timestep) as input to the modulation layers (e.g. DiT [78] modulation) which modulate the features of all layers in the network. This leads to OOD velocities distorting all features within a forward pass. The center plot of Fig. 12 shows that UPT has the best OOD generalization among all models. Notably, the performance of GINO drastically drops for higher velocities.
Generalization to different geometries
@@
The in-distribution dataset contains randomly placed circles of varying size where the mesh is generated via an adaptive meshing algorithm. To investigate robustness to the meshing algorithm, we generate a OOD dataset with a uniform meshing algorithm. In this dataset, mesh cells are approximately uniform, i.eṫhe distance between two points is roughly the same. This is in contrast to an adaptive mesh, where regions around an object have more mesh cells and therefore the distance between two points is smaller in these regions. Additionally, we investigate generalization to different obstacle geometries by using polygons (with up to 9 edges) or triangles instead of circles. Note that even though polygons are a more "complicated" shape, they are more reminiscent of a circle than triangles. The size of the obstacle is also varied here. For simplicity, the number of objects per simulation is set to 1 in this study. Note that U-Net and FNO interpolate the mesh onto a 2D grid and therefore their distribution of input positions does not change here, only the dynamics of the simulation. For GINO and UPT, also the distribution of input position changes. The right side of Fig. 12 shows that UPT achieves the best performances on OOD meshes, despite having to adjust to a different input position distribution.
@@
Figure 12: OOD generalization study. Trained 68M parameter models are evaluated on OOD datasets with more objects, higher inflow velocities and different geometries. "ID" refers to an adaptive meshing algorithm with circles as obstacles. "UMesh" changes the meshing algorithm from adaptive to uniform. "Triangle" uses triangles instead of circles and "Polygon" uses polygon obstacles (with up to 9 edges) instead of circles. UPTs have similar OOD generalization capabilities as other models, outperforming them in all evaluations. Grey indicates OOD settings.
D.6 Lagrangian fluid mechanics
Datasets
We use the Taylor-Green vortex datasets from Toshev et al. [101] for our Lagrangian experiments, comprising two dimensions (TGV2D) and three dimensions (TGV3D). The Taylor-Green vortex problem is characterized by a distinctive initial velocity field without an external driving force, resulting in a static decrease in kinetic energy over time. The TGV2D dataset includes 2500 particles, a sequence length of 126, and is characterized by a Reynolds number (Re) of 100. The data is split into 100/50/50 for training, validation, and testing, respectively. The TGV3D dataset consists of 8000 particles, a sequence length of 61, and a Reynolds number of 50. The data is partitioned into 200/100/100 for training, validation, and testing, respectively.
Baselines
The Graph Network-based Simulator (GNS) model [87] is a popular learned surrogate for physical particle-based simulations. The architecture is kept simple, adhering to the encoder-processor-decoder principle [6], where the processor consists of multiple graph network blocks [5]. The Steerable E(3)-equivariant Graph Neural Network (SEGNN) architecture [15] is a general implementation of an E($3$) equivariant GNN, where layers are directly conditioned on steerable attributes for both nodes and edges. The main building blocks are steerable MLPs, i.e., stacks of learnable linear Clebsch-Gordan tensor products interleaved with gated non-linearities [109]. SEGNN layers are message-passing layers [28] where steerable MLPs replace the traditional non-equivariant MLPs for both message and node update functions.
We utilize the checkpoints provided by Toshev et al. [101] for the comparisons in 4.3. Both GNS and SEGNN baseline models comprise 10 layers with latent dimensions of 128 and 64, respectively. The maximum irreps order of SEGNN are $l=1$, for more details see [101].
Hyperparameters
We train using a batchsize of 128 for 50 epochs with a warmup phase of 10 epochs. We sample inputs between $50\%$ to $100\%$ of total particles. For optimization, we use AdamW [59] for all experiments with learning rate $10^{-3}$ and weight decay of $0.05$.
UPT architecture
For the TGV2D and TGV3D experiments, we use $n_{\text{s}}=256$ and $n_{\text{s}}=512$ supernodes, respectively. For TGV3D more supernodes are required due to the increased number of particles. Message passing is done with a maximum of four input points per supernode. The flexibility of the UPT encoder allows us to randomly sample $50\%$ up to $100\%$ of the total particles, diversifying the training procedure and forcing the model to learn the underlying dynamics.
Experiments on Taylor Green Vortex 2D
We show results on the TGV2D dataset from LagrangeBench [101] and the time to generate a simulated trajectory in Figure 14. The velocity error behaves similar to the 3D version of the dataset in Figure 7. UPT is roughly 10 times faster than GNNs and 100 times faster than the classical SPH solver.
Model Time Speedup Device
SPH 4.9s 1x A6000
SEGNN 2.76s 2x A100
GNS 0.56s 9x A100
UPT 0.05s 98x A100
Figure 14: Left: Mean Euclidean norm of the velocity error over all particles for different timesteps. UPTs effectively learn the underlying field dynamics, resulting in lower velocity error as the trajectory evolves in time. Right: Comparison of simulation/rollout runtimes for a TGV2D trajectory with 125 timesteps and 2500 particles across SPH simulation.
D.7 Qualitative study of scaling limits details